OpenStudy (math2400):

can someone explain this derivative ? i have the answer but the steps are unclear to me

OpenStudy (math2400):

\[\frac{ d }{ dx } [\int\limits_{0}^{x^5} \cos(3t+1) dt ]\]

OpenStudy (amistre64):

what is the fundamental thrm of calculus for integration?

OpenStudy (math2400):

\[ \int\limits_a^b f(x)dx=F(b)-F(a)\]

OpenStudy (amistre64):

good, but lets use different notation, f= f', and F = f this helps to see the relationship between F and f in a more normal fashion. \[\int_{a}^{b}f'(x)~dx=f(b)-f(a)\] same setup, just more familiar notations what happens when we take the derivative of both sides?

OpenStudy (math2400):

okay got that, and um you get the derivative..? lol

OpenStudy (amistre64):

whats it look like tho, we are working a generalization

OpenStudy (math2400):

don't u have to undo the antiderivative first on the Right side?

OpenStudy (math2400):

*left

OpenStudy (amistre64):

we already have, its the right side

OpenStudy (amistre64):

the right side is the 'undoing' of the derivative of the left side. f' comes from f

OpenStudy (math2400):

oh okay so if i take the derivative of f(b)-f(a)?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

almost, a and b need not be constants, they can be functions, so we apply the chain rule \[\frac d{dx}\left[\int_{a}^{b}f'(t)~dt\right]=\frac d{dx}f(b)-\frac d{dx}f(a)\] \[\frac d{dx}\left[\int_{a}^{b}f'(t)~dt\right]=b'~f'(b)-a'~f'(a)\]

OpenStudy (amistre64):

notice here, that we already know what f' is ... we know a and b, therefore we know a' and b'

OpenStudy (math2400):

ok so u u just need to get the derivatives of a and b and multiy by f'

OpenStudy (amistre64):

\[\frac{ d }{ dx } [\int\limits_{0}^{x^5} \cos(3t+1) dt ]\] f'(a) = cos(3(0)+1) f'(b) = cos(3x^5+1) a(x) = 0, a' = 0 b(x) = x^5, b' = 5x^4

OpenStudy (amistre64):

you have all the parts required to determine the results, without havig to intgrate first, and then take the derivative ... its a shortcut if you will

OpenStudy (math2400):

so when i re-write it into the equ. i use f'(a) and f'(b) instead of putting the original....so there should be no "t" variable

OpenStudy (amistre64):

correct, the a and b are in the domain of t in this case.

OpenStudy (math2400):

i got it actually!

OpenStudy (math2400):

thank u so much :)

OpenStudy (amistre64):

youre welcome